# A Survey on Speech Large Language Models for Understanding

### v5

@@ Abstract
Index Terms:
Large Language Models, Speech Understanding
@@ I Introduction
Speech contains a wealth of useful information, ranging from linguistic content, such as lexical meaning and syntax, to paralinguistic cues like emotion and speaking style, as well as non-linguistic context such as background noise and spatial cues. Understanding speech is therefore an essential task, enabling effective human-machine communication and driving advancements in diverse fields such as virtual assistants, automatic transcription, sentiment analysis, and conversational AI[1, 2]. The research on speech understanding has a long and evolving history, deeply rooted in the intersection of speech processing and natural language processing (NLP). Its development has spanned decades, transitioning from early rule-based and statistical approaches to modern deep learning and large language model (LLM) paradigms[3].
Although the term “speech understanding” is widely used, there is currently no universally accepted definition. In this survey, we adopt a broad, task-agnostic definition: speech understanding can be defined as the integrated process that involves the perception and cognition of spoken language, ultimately producing a textual interpretation. Unlike traditional natural language understanding (NLU), which operates solely on symbolic textual input, speech understanding involves the multimodal interpretation of acoustic signals—including lexical content, prosody, speaker traits, and environmental context. It requires not only recognizing what is said, but also perceiving how it is said, who is speaking, and under what conditions. This process entails extracting diverse information types from speech, aligning them with cognitive goals, and producing structured or generative outputs, thereby bridging low-level signal processing with high-level language understanding.
This contrasts with spoken language understanding (SLU), which focuses more narrowly on extracting structured semantics from transcribed speech [4]. In comparison, our definition reflects a more holistic view that treats speech as a multimodal signal for broad understanding and reasoning tasks. In this introduction, we will start with the historical evolution of the speech understanding approaches, and subsequently introduce the primary focus of this survey: Speech Large Language Models (Speech LLMs) for understanding.
@@
The first major stage in the development of unified speech language models emerged in the late 2010s, characterized by end-to-end systems that map speech directly to high-level semantic representations, such as textual labels, summaries, or task-specific outputs [15, 5, 18, 19]. These models typically employed either encoder architectures trained from scratch using ASR modeling framework, such as CTC, RNN-T, or attention-based encoder-decoder models [20, 21, 22, 8], or self-supervised pretrained encoders such as wav2vec 2.0, HuBERT, and WavLM [23, 24, 25]. In these cases, encoders are paired with lightweight decoders to produce task-specific semantic outputs. Compared to cascaded pipelines, these models offer reduced error propagation, improved robustness, and greater architectural simplicity. However, the capacity for complex reasoning and task adaptation remains limited, primarily due to the simple decoders and the lack of general linguistic knowledge. These limitations motivate the next-stage LLM-driven approaches.
@@
Speech LLMs transition speech understanding from the encoder-decoder architectures to language-model-centric frameworks. Leveraging powerful pretrained text-based LLMs, Speech LLMs map speech inputs, typically processed by self-supervised encoders, into the textual embedding space, enabling language models to perform reasoning, text generation, and understanding directly from speech. This architecture enables a wide range of downstream tasks, including summarization, question answering, and emotion detection through flexible prompting and instruction tuning. Representative models such as Salmonn [26], SEED [27], Listen-Think-Understand [28], and DESTA-2 [29] exemplify this paradigm [30, 31, 32, 33]. More importantly, Speech LLMs integrate transcription, semantic interpretation, and response generation into a unified framework, eliminating the barrier between ASR and SLU and enabling holistic speech understanding. A detailed discussion of general speech understanding is provided in Section III.
@@
Despite the rapid progress of LLM-centric frameworks, the definition of Speech LLMs remains non-standardized in current research. Although interpretations and research scopes vary, this survey defines Speech LLMs as understanding-oriented spoken language processing. Specifically, we focus on Speech LLMs that generate textual outputs, such as transcriptions, intent labels, summaries, or answers. Alternative definitions may include speech-to-speech generation tasks, which are beyond the scope of this work and are reviewed in other surveys [34, 35]. While several recent surveys have explored the related concepts [36, 34], none have provided a focused and systematic review from the perspective of speech understanding. Existing surveys primarily focus on speech generation or broad multi-modal frameworks, offering limited context for linking Speech LLMs to traditional speech understanding tasks. This survey fills this gap by providing a focused, structured overview of the design, training, and evaluation of Speech LLMs for understanding-oriented tasks.
@@
  • We review the evolution of modeling approaches—ranging from cascaded pipelines to LLM-centric architectures—and provide a structured synthesis of current model designs, training strategies, datasets, along with an innovative evaluation of their generalization performance across speech understanding tasks.
@@
By centering the discussion on understanding-oriented tasks and offering both conceptual clarity and technical depth, this work serves as a foundational reference helping researchers in speech processing, spoken language understanding, multimodal modeling, and conversational AI better understand the current landscape and situate new models within it.
The remainder of this survey is organized as follows. Section II introduces a taxonomy of speech understanding tasks, defining the scope of speech understanding and categorizing the major task types. Section III traces the evolution of Speech LLMs, covering the shift from cascaded pipelines to unified models and contrasting architectural paradigms based on discrete versus continuous speech modeling. Section IV examines the model structure of Speech LLMs, breaking it down into three stages: modality feature extraction, modality information fusion, and LLM inference. Section V reviews training strategies of current Speech LLMs, while Section VI summarizes key datasets used for pretraining and evaluation. Section VII discusses performance benchmarks and empirical results across various understanding tasks. Finally, Section VIII outlines open challenges and future directions for advancing deep speech understanding with LLMs. The overall organization is illustrated as Figure 1.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, rounded corners, align=left, minimum width=2.5em, inner xsep=4pt, inner ysep=2pt, , where level=0align=center,font=, where level=1text width=12.8em,align=center,font=, where level=2text width=20.5em,align=left,font=, [Speech LLMs for Understanding [Sec. II: Taxonomy [Sec. II-A: Informational Dimension] [Sec. II-B: Functional Dimension] [Sec. II-C: Format Dimension] ] [Sec. III: Evolution [Sec. III-A: Emergence of Speech LLM] [Sec. III-B: Structure Evolution of Speech LLM] ] [Sec. IV: Model Structure [Sec. IV-A: Modality Feature Extraction] [Sec. IV-B: Modality Information Fusion] [Sec. IV-C: LLM Inference] ] [Sec. V: Training Strategies [Sec. V-A: Modality Alignment] [Sec. V-B: Multitask Training] [Sec. V-C: Preference Alignment] [Sec. V-D:Training of Speech LLMs with Discrete Tokens] ] [Sec. VI: Datasests [Sec. VI-A: Datasets for Perception & Shallow Cognition Tasks] [Sec. VI-B: Datasets for Deep Cognition Tasks] [Sec. VI-C: Recent Developments in Datasets] ] [Sec. VII: Performance in SU Tasks [Sec. VII-A: Taxonomy of Evaluation Methods] [Sec. VII-B: Performance of Speech LLMs Across Tasks] ] [Sec. VIII: Challenges [Sec. VIII-A: Instruction Sensitivity] [Sec. VIII-B: Degradation on Semantic Understanding Ability] ] [Sec. IX: Future Exploration] [Sec. X: Conclusion] ]
@@ II A Taxonomy of Speech Understanding Tasks
To further structure the concept of speech understanding, we categorize speech understanding tasks from a system-oriented perspective comprising three complementary dimensions. The informational dimension focuses on the types of information embedded in speech input, forming the perceptual front-end. The functional dimension considers the understanding objectives, defining what cognition tasks the system is expected to perform. Finally, the format dimension addresses how these tasks are instantiated in terms of input–output structure, capturing the overall formulation and operationalization of speech understanding systems.
@@ II-A Informational Dimension: Categorization by Speech Information Types
  • Linguistic Information: Focused on transcribing and understanding lexical and syntactic content from speech. Key tasks include Automatic Speech Recognition (ASR) and Spoken Language Translation (SLT), where challenges arise from homophones, dialectal variations, and transcription errors.
  • Paralinguistic Information: Concerned with extracting cues beyond words, such as emotion, prosody, and speaker intent. Related tasks include Emotion Recognition, Speaker Diarization and Dialogue Act Classification, which rely on features like intonation and speaking style.
  • Non-Linguistic Information: Encompasses contextual acoustic cues supporting interaction analysis. This includes tasks like Voice Activity Detection (VAD), Sound Event Classification and Acoustic Scene Understanding, which depend on background noise, speaker turns, and spatial cues.
This dimension informs the design of front-end speech encoders by highlighting the diverse information types embedded in speech. By distinguishing between linguistic, paralinguistic and non-linguistic content, it enables targeted modeling strategies for feature extraction, representation learning, and input conditioning in speech understanding systems.
@@ II-B Functional Dimension: Categorization by Cognitive Objectives
  • Perception Tasks: These focus on extracting and manipulating audio signals, requiring minimal semantic reasoning. Representative tasks include automatic speech recognition (ASR), keyword spotting (KWS), speaker diarization (SD), and voice activity detection (VAD).
  • Shallow Cognition Tasks: These rely on intuitive, pattern-based understanding, corresponding to fast and heuristic reasoning. Examples include speech translation (ST), emotion classification, intent detection, and sentiment recognition.
  • Deep Cognition Tasks: These require complex, context-aware reasoning, often involving discourse modeling and inference. Typical tasks include complex spoken question answering, conversational summarization, speaker intent explanation in ambiguous contexts, and multimodal reasoning grounded in acoustic and textual context.
Our definition of the distinction between shallow and deep cognition tasks is grounded in the Dual Process Theory of cognition, which posits two complementary modes of reasoning: System 1, which operates quickly, automatically, and intuitively; and System 2, which engages in slower, more deliberate, and analytical thinking [37, 38]. In this framework, shallow cognition tasks in speech understanding reflect System 1 processes, as they often rely on heuristic or pattern-based interpretation without the need for explicit multi-step reasoning. In contrast, deep cognition tasks correspond to System 2, requiring models to perform structured inference, integrate complex contextual cues, and resolve ambiguity. Framing speech understanding in this dual-process perspective helps offer a principled foundation for stratifying task complexity and guiding model design. By clarifying the cognitive depth required by each task type—from perceptual to deep inference—this dimension facilitates strategic planning of training objectives, curriculum learning stages, and evaluation protocols across varying levels of semantic complexity.
@@ II-C1 Input Formulation
  • Structured Input Sources: explicitly formatted signals that provide direct guidance to the model. Typical examples include:
@@
    • Lexicons or domain constraints (e.g., task-specific vocabularies or grammar rules)
    • Knowledge structures (e.g., ontologies, documents, or database schemas that encode external domain knowledge in structured form.)
  • Unstructured Input Sources: open-form, contextually derived signals that are not explicitly constrained in format. These may include:
@@ II-C2 Output Formulation
  • Unstructured Tasks: produce open-ended outputs lacking standardized answers, resisting formal evaluation and depending on contextual or subjective judgment. Representative tasks include:
@@
    • Speech translation (ST) (e.g., converting speech from one language to another with diverse valid outputs, evaluated by BLEU [39] [40]),
@@
This framework not only enables a comprehensive categorization of existing tasks, but also provides guidance for the design of future speech understanding systems. Structured outputs benefit from precise metrics and strong supervision, while unstructured outputs better reflect real-world complexity and the need for generative reasoning. Similarly, integrating structured and unstructured input sources enhances model controllability, personalization, and adaptability in diverse environments.
Taken together, when situated in the context of Speech LLMs, this taxonomy not only provides a principled framework for organizing current capabilities, but also highlights the evolving demands of future systems. With increasing emphasis on multi-task generalization, instruction following, and context-aware reasoning, Speech LLMs must flexibly integrate diverse input types and generate outputs of varying structure and specificity. The taxonomy thus informs the development of more adaptive, interpretable, and goal-aligned architectures—supporting a shift from narrow ASR-centric modeling to a broader paradigm of holistic spoken language understanding.
@@ III-A From Cascade to Unified: The Emergence of Speech LLMs for Speech Understanding
We broadly divide the development of models for speech understanding into two main stages. The first is the era of modular architectures, typically characterized by cascaded structures with separately optimized components. The second is the emergence of end-to-end architectures, where the system is trained holistically to optimize performance across the entire pipeline. Within the end-to-end paradigm, a further transition has occurred—from E2E Specialized Speech Understanding System, an end-to-end system designed for a single or a small set of well-defined tasks, to E2E General Speech Understanding Systems by allowing language instructions to specify any task and supporting outputs that can capture any form of understanding. Each stage not only represents a milestone in model capacity and architecture but also demonstrates the increasing convergence between tasks, ultimately culminating in a unified modeling capability aimed at general speech understanding. An overview of this development process is illustrated in Figure 3.

### v6

@@ I Introduction
Speech contains a wealth of useful information, ranging from linguistic content, such as lexical meaning and syntax, to paralinguistic cues like emotion and speaking style, as well as non-linguistic context, such as background noise and spatial cues. Understanding speech is therefore an essential task, enabling effective human-machine communication and driving advancements in diverse fields such as virtual assistants, automatic transcription, sentiment analysis, and conversational AI [1, 2]. Research on speech understanding has a long and evolving history, deeply rooted in the intersection of speech processing and natural language processing (NLP). Its development has spanned decades, transitioning from early rule-based and statistical approaches to modern deep learning and large language model (LLM) paradigms[3].
Although the term “speech understanding” is widely used, there is currently no universally accepted definition. In this survey, we adopt a broad, task-agnostic definition: speech understanding can be defined as the integrated process that involves the perception and cognition of spoken language, ultimately producing a textual interpretation. Unlike traditional natural language understanding (NLU), which operates solely on symbolic textual input, speech understanding involves the multimodal interpretation of acoustic signals—including lexical content, prosody, speaker traits, and environmental context. It requires not only recognizing what is said but also perceiving how it is said, who is speaking, and under what conditions. This process entails extracting diverse information types from speech, aligning them with cognitive goals, and producing structured or generative outputs, thereby bridging low-level signal processing with high-level language understanding.
This contrasts with spoken language understanding (SLU), which focuses more narrowly on extracting structured semantics from transcribed speech [4]. In comparison, our definition reflects a more holistic view that treats speech as a multimodal signal for broad understanding and reasoning tasks. In this introduction, we will start with the historical evolution of speech understanding approaches, and subsequently introduce the primary focus of this survey: Speech Large Language Models (Speech LLMs) for understanding.
@@
The first major stage in the development of unified speech language models emerged in the late 2010s, characterized by end-to-end systems that map speech directly to high-level semantic representations, such as textual labels, summaries, or task-specific outputs [15, 5, 18, 19]. These models typically employed either encoder architectures trained from scratch using ASR modeling frameworks, such as CTC, RNN-T, or attention-based encoder-decoder models [20, 21, 22, 8], or self-supervised pretrained encoders such as wav2vec 2.0, HuBERT, and WavLM [23, 24, 25]. In these cases, encoders are paired with lightweight decoders to produce task-specific semantic outputs. Compared to cascaded pipelines, these models offer reduced error propagation, improved robustness, and greater architectural simplicity. However, although task-specific models have achieved relatively mature performance in individual speech understanding tasks, they remain inadequate for the broader challenge of general speech understanding, which requires the ability to process and integrate diverse types of information from speech. The capacity for complex reasoning and task adaptation of models at this stage remains limited, primarily due to the simple decoders and the lack of general linguistic knowledge. These limitations motivate the next-stage LLM-driven approaches.
@@
Speech LLMs transition speech understanding from the encoder-decoder architectures to language-model-centric frameworks. Leveraging powerful pretrained text-based LLMs, Speech LLMs map speech inputs, which are typically processed by self-supervised encoders, into the textual embedding space. This enables language models to perform reasoning, text generation, and understanding directly from speech. This architecture enables a wide range of downstream tasks, including summarization, question answering, and emotion detection through flexible prompting and instruction tuning. Representative models such as SALMONN [26], SEED [27], Listen-Think-Understand [28], and DESTA-2 [29] exemplify this paradigm [30, 31, 32, 33]. More importantly, Speech LLMs integrate transcription, semantic interpretation, and response generation into a unified framework, eliminating the barrier between ASR and SLU and enabling holistic speech understanding. A detailed discussion of general speech understanding is provided in Section III.
@@
Despite the rapid progress of LLM-centric frameworks, the definition of Speech LLMs remains non-standardized in current research. Although interpretations and research scopes vary, this survey defines Speech LLMs as understanding-oriented spoken language processing. Specifically, we focus on Speech LLMs that generate textual outputs, such as transcriptions, intent labels, summaries, or answers. Alternative definitions may include speech-to-speech generation tasks, which are beyond the scope of this work and are reviewed in other surveys [34, 35]. While several recent surveys have explored the related concepts [36, 34], none have provided a focused and systematic review from the perspective of speech understanding. Beyond serving as a core module in most mainstream Speech LLMs, speech understanding itself has emerged as a distinct research direction, with dedicated Speech LLMs now being developed solely for understanding tasks—a trend gaining significant traction in both academia and industry[32, 28, 26, 37, 38, 39]. Existing surveys primarily focus on speech generation or broad multimodal frameworks, offering limited context for linking Speech LLMs to traditional speech understanding tasks. This survey fills this gap by providing a focused, structured overview of the design, training, and evaluation of Speech LLMs for understanding-oriented tasks.
@@
  • We review the evolution of modeling approaches—ranging from cascaded pipelines to LLM-centric architectures—and provide a structured synthesis of current model designs, training strategies, and datasets, along with an innovative evaluation of their generalization performance across speech understanding tasks.
@@
By centering the discussion on understanding-oriented tasks and offering both conceptual clarity and technical depth, this work serves as a foundational reference that helps researchers in speech processing, spoken language understanding, multimodal modeling, and conversational AI better understand the current landscape and situate new models within it.
The remainder of this survey is organized as follows. Section II introduces a taxonomy of speech understanding tasks, defining the scope of speech understanding and categorizing the major task types. Section III traces the evolution of Speech LLMs, covering the shift from cascaded pipelines to unified models and contrasting architectural paradigms based on discrete versus continuous speech modeling. Section IV examines the model structure of Speech LLMs, breaking it down into three stages: modality feature extraction, modality information fusion, and LLM inference. Section V reviews the training strategies of current Speech LLMs, while Section VI summarizes the key datasets used for pretraining and evaluation. Section VII discusses performance benchmarks and empirical results across various understanding tasks. Finally, Section VIII outlines open challenges and future directions for advancing deep speech understanding with LLMs. The overall organization is illustrated in Figure 1.
Speech LLMs for Understanding
Sec. II: Taxonomy Sec. II-A: Informational Dimension Sec. II-B: Functional Dimension Sec. II-C: Format Dimension Sec. III: Evolution Sec. III-A: Emergence of Speech LLM Sec. III-B: Structure Evolution of Speech LLM Sec. IV: Model Structure Sec. IV-A: Modality Feature Extraction Sec. IV-B: Modality Information Fusion Sec. IV-C: LLM Inference Sec. V: Training Strategies Sec. V-A: Modality Alignment Sec. V-B: Multitask Training Sec. V-C: Preference Alignment Sec. V-D: Training of Speech LLMs with Discrete Tokens Sec. VI: Datasets Sec. VI-A: Datasets for Perception & Shallow Cognition Tasks Sec. VI-B: Datasets for Deep Cognition Tasks Sec. VI-C: Recent Developments in Datasets Sec. VII: Performance in SU Tasks Sec. VII-A: Taxonomy of Evaluation Methods Sec. VII-B: Performance of Speech LLMs Across Tasks Sec. VII-C: Model Performance in Real-World Scenarios Sec. VIII: Challenges Sec. VIII-A: Instruction Sensitivity Sec. VIII-B: Degradation on Semantic Understanding Ability Sec. IX: Future Exploration Sec. X: Conclusion
@@ II A Taxonomy of Speech Understanding Tasks
To further structure the concept of speech understanding, we categorize speech understanding tasks from a system-oriented perspective, comprising three complementary dimensions. The informational dimension focuses on the types of information embedded in speech input, forming the perceptual front end. The functional dimension considers the understanding objectives, defining what cognition tasks the system is expected to perform. Finally, the format dimension addresses how these tasks are instantiated in terms of input–output structure, capturing the overall formulation and operationalization of speech understanding systems.
@@ II-A Informational Dimension: Categorization by Speech Information Types
  • Linguistic Information: Focuses on transcribing and understanding lexical and syntactic content from speech. Key tasks include Automatic Speech Recognition (ASR) and Spoken Language Translation (SLT), where challenges arise from homophones, dialectal variations, and transcription errors.
  • Paralinguistic Information: Concerned with extracting cues beyond words, such as emotion, prosody, and speaker intent. Related tasks include Emotion Recognition, Speaker Diarization, and Dialogue Act Classification, which rely on features like intonation and speaking style.
  • Non-Linguistic Information: Encompasses contextual acoustic cues supporting interaction analysis. This includes tasks like Voice Activity Detection (VAD), Sound Event Classification, and Acoustic Scene Understanding, which depend on background noise, speaker turns, and spatial cues.
This dimension informs the design of front-end speech encoders by highlighting the diverse information types embedded in speech. By distinguishing between linguistic, paralinguistic, and non-linguistic content, it enables targeted modeling strategies for feature extraction, representation learning, and input conditioning in speech understanding systems.
@@ II-B Functional Dimension: Categorization by Cognitive Objectives
  • Perception Tasks: Focus on extracting and manipulating audio signals, requiring minimal semantic reasoning. Representative tasks include automatic speech recognition (ASR), keyword spotting (KWS), speaker diarization (SD), and voice activity detection (VAD).
  • Shallow Cognition Tasks: Rely on intuitive, pattern-based understanding, corresponding to fast and heuristic reasoning. Examples include speech translation (ST), emotion classification, intent detection, and sentiment recognition.
  • Deep Cognition Tasks: Require complex, context-aware reasoning, often involving discourse modeling and inference. Typical tasks include complex spoken question answering, conversational summarization, speaker intent explanation in ambiguous contexts, and multimodal reasoning grounded in acoustic and textual context.
Our definition of the distinction between shallow and deep cognition tasks is grounded in the Dual Process Theory of cognition, which posits two complementary modes of reasoning: System 1, which operates quickly, automatically, and intuitively, and System 2, which engages in slower, more deliberate, and analytical thinking [40, 41]. In this framework, shallow cognition tasks in speech understanding reflect System 1 processes, as they often rely on heuristic or pattern-based interpretation without the need for explicit multi-step reasoning. In contrast, deep cognition tasks correspond to System 2, requiring models to perform structured inference, integrate complex contextual cues, and resolve ambiguity. Framing speech understanding in this dual-process perspective helps provide a principled foundation for stratifying task complexity and guiding model design. By clarifying the cognitive depth required by each task type—from perceptual to deep inference—this dimension facilitates strategic planning of training objectives, curriculum learning stages, and evaluation protocols across varying levels of semantic complexity.
@@ II-C Format Dimension: Categorization by Input–Output Structures
Unlike the information and functional dimensions, which describe what knowledge is used and why a task is performed, the format dimension specifies how inputs are packaged and constrained. Because structure is observable and easy to label, this dimension offers a crisp taxonomy without diluting the main theme. It also aligns with our evaluation suite in Section VII-A.
@@ II-C1 Input Formulation
  • Structured Input Sources: Provide explicitly formatted signals that provide direct guidance to the model. Typical examples include:
@@
    • Lexicons or domain constraints (e.g., task-specific vocabularies or grammar rules),
    • Knowledge structures (e.g., ontologies, documents, or database schemas that encode external domain knowledge in structured form)
  • Unstructured Input Sources: Provide open-form, contextually derived signals that are not explicitly constrained in format. These may include:
@@ II-C2 Output Formulation
  • Unstructured Tasks: Produce open-ended outputs without standardized answers, resisting formal evaluation and depending on contextual or subjective judgment. Representative tasks include:
@@
    • Speech translation (ST) (e.g., converting speech from one language to another with diverse valid outputs, evaluated by BLEU [42, 43]),
@@
This framework not only enables a comprehensive categorization of existing tasks but also provides guidance for the design of future speech understanding systems. Structured outputs benefit from precise metrics and strong supervision, while unstructured outputs better reflect real-world complexity and the need for generative reasoning. In particular, structured outputs are directly interpretable by downstream computational systems and allow reliable, objective evaluation, whereas unstructured outputs require generative understanding and are not straightforwardly machine-actionable. Similarly, integrating structured and unstructured input sources enhances model controllability, personalization, and adaptability in diverse environments.
Taken together, when situated in the context of Speech LLMs, this taxonomy not only provides a principled framework for organizing current capabilities, but also highlights the evolving demands of future systems. With increasing emphasis on multi-task generalization, instruction following, and context-aware reasoning, Speech LLMs must flexibly integrate diverse input types and generate outputs of varying structure and specificity. The taxonomy thus informs the development of more adaptive, interpretable, and goal-aligned architectures. It supports a shift from narrow ASR-centric modeling to a broader paradigm of holistic spoken language understanding.
@@ III-A From Cascade to Unified: The Emergence of Speech LLMs for Speech Understanding
We broadly divide the development of models for speech understanding into two main stages. The first is the era of modular architectures, typically characterized by cascaded structures with separately optimized components. The second is the emergence of end-to-end architectures, where the system is trained holistically to optimize performance across the entire pipeline. Within the end-to-end paradigm, a further transition has occurred. The field has moved from E2E Specialized Speech Understanding Systems, which are designed for a single or a small set of well-defined tasks, to E2E General Speech Understanding Systems. In these general systems, language instructions can specify any task, and the model supports outputs that can capture any form of understanding. Each stage not only represents a milestone in model capacity and architecture but also demonstrates the increasing convergence between tasks, ultimately culminating in a unified modeling capability aimed at general speech understanding. An overview of this development process is illustrated in Figure 3.